OpenStudy (anonymous):

How do I do this problem!!! j(j(x)) j = x-7

OpenStudy (anonymous):

is it j(x) = x-7?

OpenStudy (anonymous):

then, you would plug in x-7 for x

OpenStudy (anonymous):

so j(x-7)

OpenStudy (anonymous):

which is j(x-7) = x-14

OpenStudy (anonymous):

ok. Thanks. That's what I got..